OpenStudy (anonymous):

Line PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (campbell_st):

ok... do you know the midpoint formula?

OpenStudy (anonymous):

m=x-x/y-y ?

OpenStudy (campbell_st):

not quite there is a formula for x and y... basically its the average of 2 x values and average of 2 y values \[x_{m} = \frac{(x_{2} - x_{1})}{2}\] \[y_{m} = \frac{ (y_{2} - y_{1})}{2}\] so you have\[ x_{m} = 18.........y_{m} = -1\]

OpenStudy (anonymous):

for points with co-ordinates (x1,y1) and (x2,y2) midpoint (x,y)=( (x1+y1)/2,(x2+y2)/2)

OpenStudy (campbell_st):

so substituting to find x \[18 = \frac{25 - x_{1}}{2}\] and for y \[-1 = \frac{-5 - y_{1}}{2}\] just solve for x and y

OpenStudy (anonymous):



OpenStudy (anonymous):

sketch not to scale

OpenStudy (anonymous):

just provided for getting an idea

OpenStudy (anonymous):

use the sketch and midpoint formula to find your answer